You must evaluate a proposal to buy a new milling machine. The purchase price of the milling machine, including shipping and installation costs, is $116,000, and the equipment will be fully depreciated at the time of purchase. The machine would be sold after 3 years for $68,000. The machine would require a $7,000 increase in net operating working capital (increased inventory less increased accounts payable). There would be no effect on revenues, but pretax labor costs would decline by $44,000 per year. The marginal tax rate is 25%, and the WACC is 8%. Also, the firm spent $4,500 last year investigating the feasibility of using the machine.

  1. How should the $4,500 spent last year be handled?
    1. Last year's expenditure is considered an opportunity cost and does not represent an incremental cash flow. Hence, it should not be included in the analysis.
    2. Last year's expenditure is considered a sunk cost and does not represent an incremental cash flow. Hence, it should not be included in the analysis.
    3. The cost of research is an incremental cash flow and should be included in the analysis.
    4. Only the tax effect of the research expenses should be included in the analysis.
    5. Last year's expenditure should be treated as a terminal cash flow and dealt with at the end of the project's life. Hence, it should not be included in the initial investment outlay.

    -Select-IIIIIIIVVItem 1
  2. What is the initial investment outlay for the machine for capital budgeting purposes after the 100% bonus depreciation is considered, that is, what is the Year 0 project cash flow? Enter your answer as a positive value. Round your answer to the nearest dollar.
    $   
  3. What are the project's annual cash flows during Years 1, 2, and 3? Do not round intermediate calculations. Round your answers to the nearest dollar.
    Year 1: $   
    Year 2: $   
    Year 3: $   
  4. Should the machine be purchased?
    -Select-Yes No
Answer #1

Solution

a) Option iv is the correct, Only the tax effect of the research expenses should be included in the analysis.

b) Initial outlay of investment ; $116,000

Less: Depreciation full 100%: $116,000

Net Investment pre tax : nil [ $116,000-$116,000]

Less: Tax Saving on Depreciation @ 25% (116,000$*25%): $29,000

Net Investment post tax ( Net investment pre tax minus tax saving on depreciation) : -29,000$

Add : Depreciation [ Non Cash Expenses] : $116,000

Net Cash outflow of Investment [ outlay] ( Net investment post tax plus depreciation, -29,000$+$116,000); $87,000

The initial investment outlay for the machine for capital budgeting purposes after the 100% bonus depreciation is considered, the Year 0 project cash flow = $87,000

c)

Present Cash Inflows
Particulars Year 1 Year 2 Year 3
Pre tax Labor Cost (in $) 44,000 44,000 44,000
Less: Tax Expenses @ 25% (in $) 11,000 11,000 11,000
a) Post tax Labour Cost (in $) 33,000 33,000 33,000
b) Less: Increase in Working Capital (in $) 7,000 7,000 7,000
c) Free Cash Flows [a-b] (in $) 26,000 26,000 26,000
d) Discounting factor post tax is 8%(1-0.25)= 6% 0.9433 0.8899 0.8396
e) Discounted Free Cash Flows (c*d) 24,525.8 23,114 21,829.60

1)Total Discounted Free Cash Flows (Total of e) $69,469.40

2)Tax Benefit on Feasibility study $4,500*25%= $1125

2 a) Present Value of Tax Benefits : $1125*0.8396= $944.55

3) Tax Expense on Sale of Assets 68000*25%= $ 17,000

3 a)Present Value of Tax Expenses: 17,000*0.8396= $ 14,273.20

Net Cash Inflows

1)Total Discounted Free Cash Flows (Total of e) $69,469.40

2 a) Present Value of Tax Benefits : $1125*0.8396= $944.5

3) Cash inflow from Sales of Assets 68000*0.8396= $57,092.80

Less : Present Value of Tax Expenses: 17,000*0.8396= $ 14,273.20

4)Present Value of Sale of Assets at year 3 [$ 57,092.80-$14,273.20]= $42,819.60

5) Net Cash Inflows ( 1+2+4) = $113,233.50

d) Net Present Value = Net Cash Inflows - Net Cash outflows [$113,233.50- $87,000]= $26,233.50

The machine should be purchased as it gives net present value in positive , as inflows exceeds outflows considering the tax and discount rates.

Note: Feasibility Study is sunk cost and not included in outlay , only its tax effect will effect the cash flows.